How to solve:
1/4 (w+2) = 1/6 (5-w)

Answer by rfer(16322)

1/4w+2/4=5/6-1/6w
1/4w+1/6w=5/6-2/4
3/12w+2/12w=10/12-6/12
5/12w=4/12
w=(4/12)/(5/12)
w=4/12*12/5
w=48/60
w=4/5